# ModalTune: Fine-Tuning Slide-Level Foundation Models with Multi-Modal Information for Multi-task Learning in Digital Pathology

### v1

@@ 11 Ablation Studies
In this section, we investigate multiple key design choices in our study: the effect of modality adapters, the effect of text embeddings, the impact of training solely on general prompts, the impact of different text encoders and projectors, illustrated in Tab. 8.
11.1 Modality Adapters
To assess more deeply if the Modal Adapter architecture provides benefits in uni-modal fine-tuning, we evaluate the performance of ModalTune by replacing transcriptomics tokens from the genomic encoder with the same number and dimension of randomly initialized trainable embedding vectors (‘Single Modal’). We find that overall, the model outperforms all other image-only models (Tab. 2) in subtype classification and has competitive performance in survival prediction. This effect is most pronounced when compared against the Gigapath fully fine-tuned model, where the model demonstrates superior performance across cancer subtype classifications and survival prediction (0.9%; 3.4%). Fine-tuning with the Modal Adapter setup requires updating fewer parameters than fully tuning Gigapath. Thus, this experiment demonstrates both the efficiency and effectiveness of the proposed architecture.
@@ 11.2 Multi-task using Texts
To examine the effect of multi-task learning, we train the Modal Adapter in a single-task manner, without embedding the tasks using text (‘No Text Embedding’). We found cancer subtype prediction (0.6% drop) was less affected than survival prediction (2.7% drop), indicating the latter utilized more information from other tasks than the former. These findings indicate the utility of using text embeddings for multi-task learning and suggest that inter-task information is beneficial for downstream performance. We additionally conduct experiments where task-specific texts are used instead of directly training on labels and observe similar performance to the No Text Embedding setup.
@@ 11.5 Projectors
While unorthodox, these findings do align with prior studies highlighting the utility of random fixed projectors in extracting non-trivial features in various scenarios. Of particular relevance to ModalTune, random projectors are effective feature extractors, reducing dimensionality and producing powerful representations [76, 3, 55]. Additionally, random projectors largely preserve inter-sample distances, as discussed in [24, 76], i.e., they maintain smaller distances between samples of the same class and larger distances between samples from different classes. This is evident empirically through the performance of linear regression on text embeddings (Tab. 7) with and without random projectors and theoretically from the Johnson-Lindenstrauss lemma [35], as discussed in Boutsidis et al. [7]. Given that clusters in the dataset are largely preserved and we opt to perform simple linear probing on extracted features, we expect the fixed random projector to be a viable, generalizable, and effective projection method for ModalTune.

### v2

@@ 11 Ablation Studies
In this section, we investigate multiple key design choices in our study: impact of different Modal Adapters, the effect of Modal Adapters, the effect of text embeddings, the impact of training solely on general prompts, the impact of different text encoders and Projectors, illustrated in Tab. 8.
11.1 Modal Adapters
Different Modal Adapters (LoRA): We compared ModalTune by extending LoRA [26], instead of using the ViT Adapter–based Modal Adapter, to handle transcriptomics and its interactions with the slide encoder. Overall, we observed that LoRA slightly underperformed ModalTune in both cancer subtype classification (1.5% drop) and survival prediction (0.8% drop), thereby motivating our choice of the Modal Adapter architecture.
Effect of Modal Adapters: To assess more deeply if the Modal Adapter architecture provides benefits in uni-modal fine-tuning, we evaluate the performance of ModalTune by replacing transcriptomics tokens from the genomic encoder with the same number and dimension of randomly initialized trainable embedding vectors (‘Single Modal’). We find that overall, the model outperforms all other image-only models (Tab. 2) in subtype classification and has competitive performance in survival prediction. This effect is most pronounced when compared against the Gigapath fully fine-tuned model, where the model demonstrates superior performance across cancer subtype classifications and survival prediction (0.9%; 3.4%). Fine-tuning with the Modal Adapter setup requires updating fewer parameters than fully tuning Gigapath. Thus, this experiment demonstrates both the efficiency and effectiveness of the proposed architecture.
@@ 11.2 Multi-task using Texts
To examine the effect of multi-task learning, we train the Modal Adapter in a single-task manner, without embedding the tasks using text (‘No Text Embedding’). We found cancer subtype prediction (0.6% drop) was less affected than survival prediction (2.7% drop), indicating the latter utilized more information from other tasks than the former. These findings indicate the utility of using text embeddings for multi-task learning and suggest that inter-task information is beneficial for downstream performance.
We also tested the utility of text embeddings for multi-task learning on other architectures such as ABMIL (cat) (‘ABMIL (cat) w/ text emb.’), and found that compared to ABMIL (cat) trained on single tasks, there was an overall drop in performance for cancer subtype classification (1.6% drop), while performance for survival prediction remained similar. However, better architectures like ModalTune, incorporating interaction terms, were able to achieve substantially improved performance when trained with text embeddings.
@@ 11.5 Projectors
While unorthodox, these findings do align with prior studies highlighting the utility of randomly-initialized and fixed projectors in extracting non-trivial features in various scenarios. Of particular relevance to ModalTune, random projectors are effective feature extractors, reducing dimensionality and producing powerful representations [76, 2, 55]. Additionally, random projectors largely preserve inter-sample distances, as discussed in [22, 76], i.e., they maintain smaller distances between samples of the same class and larger distances between samples from different classes. This is evident empirically through the performance of linear regression on text embeddings (Tab. 7) with and without random projectors and theoretically from the Johnson-Lindenstrauss lemma [34], as discussed in Boutsidis et al. [6]. Given that clusters in the dataset are largely preserved and we opt to perform simple linear probing on extracted features, we expect the fixed random Projector to be a viable, generalizable, and effective projection method for ModalTune.
@@
12 Experiments with TITAN
To demonstrate the ability of ModalTune to extend to other Transformer-based SLFMs, we perform experiments interfacing it with the TITAN [17] SLFM. To do so, we first re-extract patch features using the method described in the original work. We then perform analogous comparisons to Gigapath-ModalTune (Tab. 9). We found TITAN to be a much stronger standalone model than Gigapath, obtaining strong results with only linear probing or fine-tuning. Nonetheless, we find TITAN benefits from the bulk transcriptomics modality over uni-modal fine-tuning (1.3% in subtype classification, 1.0% in survival prediction), and slightly improves when tuned using the ModalTune pipeline (1.1%, 0.4%).
We additionally probe the generalizability of ModalTune TITAN similarly to the generalization study in Sec. 4.4 (Tab. 11). We found the ModalTune framework to greatly benefit OOD prediction performance on COADREAD and BLCA, with an average of 13.9% improvement in subtype prediction and 2.9% improvement in survival prediction than the next best OOD baseline. Furthermore, ModalTune demonstrated generalizability, performing only 0.4% worse than a fully-supervised TITAN (cat) network in subtype prediction and 6.3% worse in survival prediction. Thus, even though we obtained modest improvements from in-domain validation, we emphasize ModalTune maintains SLFM generalization better than conventional tuning methods.
13 Additional Modalities
To validate ModalTune’s extensibility, we integrated clinical data () for TCGA by incorporating available features: patient age (all); TNM staging (BRCA, NSCLC, RCC); treatment type (BRCA only); and hormone receptor status (BRCA only). A 2-layer MLP encodes into for concatenation with transcriptomics [Sec. 3.2]. Overall, in our experiments shown in Tab. 10, the addition of clinical data marginally improved ModalTune’s performance, primarily for survival prediction, while in the case of RCC subtype classification, it even led to a degradation, possibly because the clinical features for RCC are less relevant for the subtype classification task. This highlights ModalTune’s ability to integrate salient features across modalities, though incorporating additional modalities remains a direction for future work.